OpenStudy (anonymous):

Pentagon RSTUV is circumscribed about a circle. What is the value of x if RS=8, ST=12, TU=15, UV=9, and V=12?

OpenStudy (anonymous):

Picture

OpenStudy (anonymous):

\[x + y = 8\]\[y + z = 12\]\[z + a = 15\]\[a + b = 9\]\[b + x = 12\] \[x - z = -4\] \[x + a = 11\] \[x - b = 2\] \[x + x = 14\]\[2x = 14\]\[x = ?\]

OpenStudy (anonymous):

I basically used the elimination method to continuously cancel the extra variables out. I also used the congruent tangents theorem.

OpenStudy (anonymous):

Thank you. 2x=14 would turn into 7 correct?

OpenStudy (anonymous):

Yes :)